OpenStudy (anonymous):

2i (7−4i) how do you distribute this??

OpenStudy (anonymous):



OpenStudy (mathteacher1729):

You would distribute it as you would for 2*(7-4) EXCEPT that i times i = -1.

OpenStudy (anonymous):

i know the aswer is 8+14i but how does that changes from 14i-8i ?

OpenStudy (mathteacher1729):

2i (7−4i) 2i*7 - 2i*4i ===> I distributed the 2i to both sides 14i - 8*i*i 14i -8*(-1) ===> because i*i = -1 14i +8 ===> because -1 times -1 = +1 And that's the steps. :)

OpenStudy (anonymous):

thank you :)

OpenStudy (mathteacher1729):

You're welcome!